OpenStudy (tootsi123):

The slope of the line whose equation is x - 3y = 1 is

OpenStudy (notcheatinglol):

What is your question?

OpenStudy (tootsi123):

A. -3 B. -1/3 C. 1/3

OpenStudy (czesc):

To find the slope simply solve for y.

OpenStudy (tootsi123):

Okay solving for y so then 3y=1-x

OpenStudy (welshfella):

rearrange to the form y = mx + b the slope is the value of m

OpenStudy (tootsi123):

correct so far

OpenStudy (czesc):

yes

OpenStudy (czesc):

get rid of the 3 on the left side.

OpenStudy (tootsi123):

divide both sides by 3?

OpenStudy (czesc):

yes

OpenStudy (welshfella):

3y = x - 1 now divide throught by 3

OpenStudy (tootsi123):

So the answer should be 1/3 is the slope correct?

OpenStudy (welshfella):

yes

OpenStudy (tootsi123):

Okay :) thank you when i first did it for some reason i kept getting 3 but i know what i did wrong now.

OpenStudy (welshfella):

y = (1/3) x - 1/3

OpenStudy (czesc):

Good job, see you around!

OpenStudy (tootsi123):

One more i just want to make sure i have it down

OpenStudy (tootsi123):

x+y=6 y=6+x correct so far?? @czesc

OpenStudy (welshfella):

No you need to subtract x from both sides

OpenStudy (tootsi123):

Oh okay y=6-x now what

OpenStudy (welshfella):

x + y = 6 -x -x y = 6 - x

OpenStudy (tootsi123):

Do I need to divide still or is it just one since it's only x?

OpenStudy (welshfella):

what do you need here - the slope again?

OpenStudy (tootsi123):

I need to know the slope but I want to make sure I know how to do these problems for my upcoming test.

OpenStudy (welshfella):

oh ok compare what we have with the general form which is y = mx + b we can write y = 6 - x as y = -x + 6 so comparing:- y = mx + b y = -x + 6

OpenStudy (welshfella):

so m = what?

OpenStudy (tootsi123):

Oh okay that makes since okay I get it now.

OpenStudy (welshfella):

-x can be written -1x so m = slope = -1

OpenStudy (tootsi123):

Alright that makes since thank you so much I really appreciate it

OpenStudy (welshfella):

yw